# The Uniqueness of LLaMA3-70B with Per-Channel Quantization: An Empirical Study

### v1

@@ Abstract
We have observed a distinctive quantization-related behavior in the LLaMA3/3.1-70B models that is absent in both the LLaMA2-70B and LLaMA3/3.1-8B/405B models. Quantization is a crucial technique for deploying large language models (LLMs) efficiently, as it reduces memory consumption, decreases memory transactions, and potentially accelerates inference using lower-precision compute cores. Among various bit widths and representations for weights and activations, the 8-bit integer weight and 8-bit integer activation (W8A8) configuration is particularly popular due to its widespread hardware support. However, the impact of W8A8 post-training quantization on model accuracy remains contentious. While several studies have suggested calibrating either weights or activations to mitigate accuracy degradation, a comprehensive solution has yet to be identified. In this paper, we empirically investigate multiple LLMs featured on an open LLM leaderboard, discovering that the LLaMA3-70B model series have a unique accuracy degradation behavior with W8A8 per-channel post-training quantization. In contrast, other model series such as LLaMA2, LLaMA3-8B, Qwen, Mixtral, Mistral, Phi-3, and Falcon demonstrate robust performance with W8A8, sometimes surpassing their FP16 counterparts. Contrary to previous assertions attributing degradation to the large dynamic range of activations, our findings indicate that the weight distribution of the LLaMA3-70B is the primary factor behind the vulnerability. By meticulously analyzing the distinct characteristics of weight distributions across Transformer blocks, we propose a mixed strategy with less than 3% of the layers enabling finer W8A8 quantization granularity, while the remaining 97% of layers retain the per-channel configuration. As a result, the average accuracy of LLaMA3-70B-W8A8 is increased from 45.5% to 73.4% (just 0.7% shy of LLaMA3-70B-FP16) across eight reasoning tasks. Notably, our method requires neither calibration nor fine-tuning.
@@ 1 Introduction
To elucidate the unique vulnerability of the LLaMA3-70B series to W8A8 per-channel quantization, we conducted a comparative analysis of its weight distributions against those of other robust models. Our investigation revealed a significant discrepancy, particularly pronounced in the initial layers. These layers exhibit a distinctive characteristic wherein magnitudes of certain weights exceed others by several orders of magnitude. Figure 2 shows an example of the “V” matrix in the first Transformer block of the LLaMA3-70B model. Weights with large magnitudes are concentrated on specific input dimensions, with the largest values exceeding 90. These outlier weights substantially expand the quantization range, resulting in larger quantization intervals and consequently diminished precision for smaller weight values. Notably, these weight outliers are only prominent in the “Q”, “K”, “V”, “Up”, and “Gate” weight matrices of the initial transformer blocks. To mitigate these quantization errors, we propose implementing per-group quantization specifically for these affected layers. Although these layers constitute merely 2-3% of the total layers in LLaMA3-70B, this targeted approach yields a dramatic enhancement in accuracy, elevating model performance to a level approaching that of FP16 models.
@@
To address this, we propose a mixed per-channel/per-group quantization strategy. This approach applies per-group quantization to less than 3% of the layers, specifically those with significant weight outliers, while maintaining per-channel quantization for the remaining 97% layers. This mixed strategy effectively restores the accuracy of the LLaMA3-70B model series to levels comparable to its FP16 counterparts.
@@ 2 Quantization Basics
The scale-factor plays a crucial role in the quantization process. The number of floating-point values used to determine max_abs is typically referred to as the group size. In LLMs, consider a weight matrix of dimension and an input activation matrix of dimension . The resultant output of has dimension . For per-channel quantization of , the group size is , implying that each row of values in shares a common scale-factor. Consequently, all scale-factors form a vector of length . Similarly, per-channel quantization of also has a group size of , with each column of values in sharing a scale-factor, resulting in a scale-factor vector of length . This configuration enables the computation of floating-point matrix multiplication using integer-bit compute cores (e.g., INT8 tensor cores), followed by dequantization. The dequantization process involves an element-wise multiplication with the scale-factor matrix of dimension , derived as the outer product of the scale-factor vectors of and . Notably, this dequantization can be efficiently fused into the subsequent operation after in LLMs, which is typically an element-wise operation such as rotary embedding, SiLU activation application, or normalization. In contrast, per-group quantization of and employs a finer granularity, where each row of (and column of ) is partitioned into groups, each with its own scale-factor. Consequently, the output cannot be computed in a single cycle within the INT tensor core, as the scale-factors differ during accumulation in the multiply-accumulation (MAC) unit. This leads to a degradation in the compute efficiency of the quantized model. Given these considerations, per-channel quantization is generally preferred unless the quantization-induced error degrades the model’s accuracy to an unacceptable level.
@@ 3.1 Test Dataset
We evaluated LLMs across eight reasoning tasks: Hellaswag (HS), PIQA (PQ), OpenbookQA (OQ), ARC-Easy (AE), Winogrande (WG), ARC-Challenge (AC), BoolQ (BQ), and MMLUZellers et al. (2019); Bisk et al. (2019); Mihaylov et al. (2018); Clark et al. (2018); Sakaguchi et al. (2019); Clark et al. (2019); Hendrycks et al. (2021a, b). We realize that many studies report the average accuracy over these reasoning tasks by simply computing the arithmetic mean of individual task accuracies (denoted by “AVG” in our figures). However, this approach may lack rigor due to the significant variance in the number of questions across tasks. For instance, HS and MMLU each contain over 10,000 questions, while OQ comprises only 500.
@@ 3.5 LLaMA3-70B is the ONLY ONE sensitive to W8A8
Figure 4 and Figure 5 compare the accuracy of other models with different architectures with W8A8 quantization. The analysis spans the LLaMA family, including LLaMA3-8B, LLaMA2-70B, and LLaMA2-70B-chat, as well as other prominent Transformer-based designs such as Qwen2-72B, Mixtral-8x7B, Mistral-123B, Phi3-13B, and Falcon-40B. These architectures represent the top-performing architectures in the LLM Open Leaderboard.
The data presented in these figures illustrates a contrast to the behavior observed in the LLaMA3-70B series. Specifically, the models examined here exhibit remarkable resilience to W8A8 per-channel quantization. In the majority of instances, the performance of W8A8 quantized versions closely approximates that of their FP16 counterparts. While minor variations exist, the performance decline rarely exceeds 1.5% (Mistral-123B).
@@ 3.6.1 A qualitative exploration
Figures 2, 6, 7, and 8, illustrate the absolute values of the V matrix weights in the first Transformer block for LLaMA3-70B, LLaMA3.1-70B, LLaMA3-8B, and LLaMA2-70B models, respectively. Other weight matrices for other models are shown in the Appendix. It is important to note that in per-channel quantization, each group comprises a length- vector along the input dimension. Our analysis reveals several key observations:
@@ 3.6.2 A quantitative exploration
Figure 9, 10, 11, 12 illustrate the max_abs and quantization error for four LLaMA models: LLaMA3-70B, LLaMA2-70B, LLaMA3-8B, and the recently released LLaMA3.1-405B. The x-axis represents the total number of layers, excluding the initial embedding layer and the final output head. Each Transformer block in a LLaMA model comprises seven matrices (Q, K, V, O, up, gate, down), with 80, 80, 32, and 126 blocks in the four models, respectively. The left y-axis denotes the quantization error, while the right y-axis indicates the max_abs weight of each layer.
@@ 4 A Mixed Grouping Strategy
Figure 13 presents a comparative analysis of the accuracy achieved by our proposed method against that of FP16 models. The presented models includes several LLaMA3-70 and its variants and derivatives in the Open LLM Leaderboard. The results convincingly demonstrate that our mixed quantization approach—combining 2.68% per-group and 97.32% per-channel 8-bit quantization—can match the accuracy of the LLaMA3-70B model with FP16 precision.
Figure 14 illustrates the significant reduction in quantization errors achieved for the initially affected layers (specifically, the Q, K, V, Up, and Gate matrices of Blocks 0, 1, and 3) through the implementation of per-group quantization with a group size of 1024. While the maximum absolute value (max_abs) for each layer remains unchanged, the reduction in the number of quantized values per group leads to a marked decrease in quantization error.
@@
5 Ablation on the group size
@@ 7.1 Post-training quantization
PTQ does not use any back-propagation to finetune weights. For weight-only quantization, LLM.int8() Dettmers et al. (2022a) uses a mixed precision with 16-bit MAC for outliers. GPTQ Frantar et al. (2023) and AWQ Lin et al. (2024a) quantize the per-group weight-only tensor using as small as 3-4 bits. SqueezeLLM Kim et al. (2024) uses non-uniform quantization and store outliers efficiently. For weight-activation quantization that enables faster inference Dettmers et al. (2022b); Wei et al. (2023) SmoothQuant Xiao et al. (2024) balances the outliers from activation to weights to enable per-tensor W8A8. QServe Lin et al. (2024b) proposes W4A8K4 to serve LLM models. Some other studies enable even smaller number of bits Shao et al. (2024); Zhao et al. (2024); Ashkboos et al. (2024).
@@ 9.1 Models
| Mistral-L-Instruct-123B | https://huggingface.co/mistralai/Mistral-Large-Instruct-2407 |
@@ 9.2.1 The first block has more large outliers
We show the distribution of weights in the first block in LLaMA3-70B. Figures 16, 17, 18, 19, 20, 21, 22 shows seven weight matrices in the 0-th Transformer block. The z-axis shows the magnitudes of outliers in these matrices. We can see that the Q, K, V, Up, and Gate matrices have larger outliers and they form a “wall” in some input indices. The other two weights, namely O and Down, do not show such a pattern.

### v2

# The Uniqueness of LLaMA3-70B Series with Per-Channel Quantization

@@ Abstract
We have observed a distinctive quantization-related behavior in the LLaMA3/3.1-70B models that is absent in both the LLaMA2-70B and LLaMA3/3.1/3.2-1B/3B/8B/405B models. Quantization is a crucial technique for deploying large language models (LLMs) efficiently, as it reduces memory consumption, decreases memory transactions, and potentially accelerates inference using lower-precision compute cores. Among various bit widths and representations for weights and activations, the 8-bit integer weight and 8-bit integer activation (W8A8) configuration is particularly popular due to its widespread hardware support. However, the impact of W8A8 post-training quantization on model accuracy, especially on the recently released LLaMA3/3.1 model series, remains contentious. While several studies have suggested calibrating either weights or activations to mitigate accuracy degradation, a comprehensive solution has yet to be identified. In this paper, we explore three key questions: What makes the LLaMA3-70B model series uniquely vulnerable to quantization? Why is this the case? And how can the issue be addressed? We empirically investigate multiple LLMs featured on an open LLM leaderboard, discovering that the LLaMA3-70B model series have a unique accuracy degradation behavior with W8A8 per-channel post-training quantization. In contrast, other model series such as LLaMA2, LLaMA3/3.1-8B, LLaMA3.2-1B/3B, Qwen, Mixtral, Mistral, Phi-3, and Falcon demonstrate robust performance with W8A8, sometimes surpassing their FP16 counterparts. Contrary to previous assertions attributing degradation to the large dynamic range of activations, our findings indicate that the weight distribution of the LLaMA3-70B is the primary factor behind the vulnerability. By meticulously analyzing the distinct characteristics of weight distributions across Transformer blocks, we propose two solutions that make different tradeoffs in hardware/software overhead. First, we propose a mixed strategy where less than 3% of the layers employ finer per-group W8A8 quantization granularity, while the remaining 97% retain the per-channel configuration. Second, we introduce a bi-smoothing strategy that balances quantization errors between weights and activations while maintaining per-channel quantization throughout. Experimental results demonstrate that both strategies effectively preserve the accuracy of the entire LLaMA3-70B model series under W8A8 quantization, achieving performance on par with their FP16 counterparts.
@@ 1 Introduction
To elucidate the unique vulnerability of the LLaMA3-70B series to W8A8 per-channel quantization, we conducted a comparative analysis of its weight distributions against those of other robust models. Our investigation revealed a significant discrepancy, particularly pronounced in the initial layers. These layers exhibit a distinctive characteristic wherein magnitudes of certain weights exceed others by several orders of magnitude. Figure 2 shows an example of the “V” matrix in the first Transformer block of the LLaMA3-70B model. Weights with large magnitudes are concentrated on specific input dimensions, with the largest values exceeding 90. These outlier weights substantially expand the quantization range, resulting in larger quantization intervals and consequently diminished precision for smaller weight values. Notably, these weight outliers are only prominent in the “Q”, “K”, “V”, “Up”, and “Gate” weight matrices of the initial transformer blocks, while the "O" and "Down" weight matrices do not exhibit this behavior. We propose two methods to address this issue. First, to mitigate these quantization errors, we propose implementing per-group quantization specifically for these affected layers. Although these layers constitute merely 2-3% of the total layers in LLaMA3-70B, this targeted approach yields a dramatic enhancement in accuracy, elevating model performance to a level approaching that of FP16 models. Second, we demonstrate a bi-directional smoothing method of the weight and activation matrices that effectively reduces the dynamic range of both, significantly minimizing quantization errors. This technique preserves per-channel quantization while enhancing the model’s resilience to W8A8 quantization.
@@
To address this, we first propose a mixed per-channel/per-group quantization strategy. This approach applies per-group quantization to less than 3% of the layers, specifically those with significant weight outliers, while maintaining per-channel quantization for the remaining 97% layers. This mixed strategy effectively restores the accuracy of the LLaMA3-70B model series to levels comparable to its FP16 counterparts.
The second method introduces bi-directional smoothing to mitigate outliers in both weights and input activations. This approach significantly reduces the dynamic range required for quantization, enabling the LLaMA3-70B model series to retain accuracy comparable to their FP16 counterparts.
@@ 2 Quantization Basics
The scale-factor plays a crucial role in the quantization process. The number of floating-point values used to determine max_abs is typically referred to as the group size. In LLMs, consider a weight matrix of dimension and an input activation matrix of dimension . The resultant output of has dimension . Here , , and represent the sequence length, hidden dimension, and output dimension in a Transformer layer, respectively. For per-channel quantization of , the group size is , implying that each row of values in shares a common scale-factor. Consequently, all scale-factors form a vector of length . Similarly, per-channel quantization of also has a group size of , with each row of values in sharing a scale-factor, resulting in a scale-factor vector of length . This configuration enables the computation of floating-point matrix multiplication using integer-bit compute cores (e.g., INT8 tensor cores), followed by dequantization. The dequantization process involves an element-wise multiplication with the scale-factor matrix of dimension , derived as the outer product of the scale-factor vectors of and . Notably, this dequantization can often be efficiently fused into the subsequent operation after in LLMs, which is typically an element-wise operation such as rotary embedding, SiLU activation application, or normalization. In contrast, per-group quantization of and employs a finer granularity, where each row of (and of ) is partitioned into groups, each with its own scale-factor. Consequently, the output cannot be computed in a single cycle within the INT tensor core, as the scale-factors differ during accumulation in the multiply-accumulation (MAC) unit. This leads to a degradation in the compute efficiency of the quantized model. Given these considerations, per-channel quantization is generally preferred unless the quantization-induced error degrades the model’s accuracy to an unacceptable level.
@@ 3.1 Test Dataset
We evaluated LLMs across eight reasoning tasks: Hellaswag (HS), PIQA (PQ), OpenbookQA (OQ), ARC-Easy (AE), Winogrande (WG), ARC-Challenge (AC), BoolQ (BQ), and MMLUZellers et al. (2019); Bisk et al. (2019); Mihaylov et al. (2018); Clark et al. (2018); Sakaguchi et al. (2019); Clark et al. (2019); Hendrycks et al. (2021a, b). We report the “accuracy” for all tasks. Some tasks have “normalized accuracy” and the conclusions derived from “accuracy” remain consistent when applied to “normalized accuracy”. We realize that many studies report the average accuracy over these reasoning tasks by simply computing the arithmetic mean of individual task accuracies (denoted by “AVG” in our figures). However, this approach may lack rigor due to the significant variance in the number of questions across tasks. For instance, HS and MMLU each contain over 10,000 questions, while OQ comprises only 500.
@@ 3.5 LLaMA3-70B is the ONLY ONE sensitive to W8A8
Figure 4 , Figure 6 and Figure 7 compare the accuracy of other models with different architectures with W8A8 quantization. The analysis spans the LLaMA family, including LLaMA3-8B, LLaMA2-70B, and LLaMA2-70B-chat, as well as other prominent Transformer-based designs such as Qwen2-72B, Mixtral-8x7B, Mistral-123B, Phi3-13B, and Falcon-40B. These architectures represent the top-performing architectures in the LLM Open Leaderboard.
The data presented in these figures illustrates a contrast to the behavior observed in the LLaMA3-70B series. Specifically, the models examined here exhibit remarkable resilience to W8A8 per-channel quantization. In the majority of instances, the performance of W8A8 quantized versions closely approximates that of their FP16 counterparts.
@@ 3.6.1 A qualitative exploration
Figures 2, 8, 9, and 10, illustrate the absolute values of the V matrix weights in the first Transformer block for LLaMA3-70B, LLaMA3.1-70B, LLaMA3-8B, and LLaMA2-70B models, respectively. Other weight matrices for other models show similar trend and are shown in the Appendix. It is important to note that in per-channel quantization, each group comprises a length- vector along the input dimension. Our analysis reveals several key observations:
@@ 3.6.2 A quantitative exploration
Figure 11, 12, 13, 14 illustrate the max_abs and quantization error for four LLaMA models: LLaMA3-70B, LLaMA2-70B, LLaMA3-8B, and the recently released LLaMA3.1-405B. The x-axis represents the total number of layers, excluding the initial embedding layer and the final output head. Each Transformer block in a LLaMA model comprises seven matrices (Q, K, V, O, up, gate, down), with 80, 80, 32, and 126 blocks in the four models, respectively. The left y-axis denotes the quantization error, while the right y-axis indicates the max_abs weight of each layer.
@@ 4 A Mixed Grouping Strategy
4.1 Main experimental results on mixed grouping
Figure 15 and Figure 16 presents a comparative analysis of the accuracy achieved by our proposed method against that of FP16 models. The presented models includes several LLaMA3-70 and its variants and derivatives in the Open LLM Leaderboard. The results convincingly demonstrate that our mixed quantization approach—combining 2.68% per-group and 97.32% per-channel 8-bit quantization—can match the accuracy of the LLaMA3-70B model with FP16 precision.
@@ 4.1 Main experimental results on mixed grouping
4.2 Ablation on the group size
@@
5 A Bi-smoothing Strategy
In Xiao et al. (2024), it was proposed to mitigate activation outliers by reducing the maximum absolute values of activations while amplifying the maximum absolute values of weights. This approach assumes that activation outliers are concentrated in certain channels and are significantly more severe than weight outliers. However, as demonstrated in previous sections, this assumption no longer holds for the LLaMA3-70B and LLaMA3.1-70B model series, including their instruction-tuned versions. In this section, we introduce bi-directional smoothing algorithms that balance the magnitudes of both weights and activations.
5.1 Equivalence in bi-directional smoothing
Let be the weights and be the input activation. In the representation of a Transformer linear layer, is output dimension, is the hidden dimension, is the batch size and is the sequence length. The output of the linear layer is calculated by
where we use Python tensor representation for the 3D tensor , and 2D tensor .
There is an equivalence when multiplying and .
Theorem 5.1.
Let be a non-zero vector. We denote and the column-wise (last dimension) smoothed tensors, respectively, such that
and
Then the following equation holds,
Proof.
Let and , then ,
We call the vector the smooth-factor for and .
5.2 How to select the smooth-factor?
For per-channel quantization, a quantization group consists of numbers and the max_abs in the quantization group determines the scale-factor and quantization interval. First, assume the batch size is 1, and then . Since it is desirable to minimize the max_abs of both weight and input activation, but the multiplication of these two values, i.e., and , remains constant to guarantee the output of is unchanged. Therefore, we choose the smooth-factor as
Then we have
i.e., the max_abs after the bi-smoothing for weight and input activation is balanced.
In general when the batch size is greater than 1, then we can choose as
5.3 Main experiment results for bi-smoothing
We show the accuracy comparison of all four LLaMA3-70B series of models (LLaMA3.1-70B-Instruct, LLaMA3.1-70B, LLaMA3-70B-Instruct, and LLaMA3-70B) in the section.
Figure 18 demonstrates the effectiveness of the proposed bi-smoothing method for the entire LLaMA3-70B model series. It is observed that the accuracy of models with per-channel W8A8 quantization are as good as their FP16 counterparts, which shows significant improvement over W8A16 models in Figure 1. Note that to calculate the smooth-factor in Equation (1), we need some calibration data to estimate . To prevent data contamination, we use only eight questions from the GPQA reasoning task for estimation. The calibration process takes only a few seconds, requiring a single inference pass over these samples. As demonstrated in the later ablation studies, the accuracy of the bi-smoothing method remains stable across varying amounts and sources of calibration data.
5.3.1 Why bi-smoothing works well?
Figure 19 compares the weights and input activations of the first Q matrix in the LLaMA3.1-70B-Instruct model. The input activations are drawn from a sequence of length 243 in the PIQA dataset. Notably, problematic weight outliers, with magnitudes exceeding 60, are suppressed to below 2.0, alongside activation outliers. This smoothing process makes both weights and activations far more resilient to quantization, achieving W8A8 accuracy nearly identical to that of the W16A16 configuration (as detailed in the next section).
In general, Figure 20 illustrates the quantization error and max_abs across all layers of the LLaMA3.1-70B-Instruct model after bi-smoothing. Compared to Figure 11, depicting the unsmoothed model, there is a substantial reduction in max_abs, yielding an order-of-magnitude decrease in quantization error.
5.4 Ablation study on bi-smoothing
From Figure 18, we can see the best performing model is LLaMA3.1-70B-Instruct. So, we use this model for ablation study of bi-smoothing algorithms. We will show the bi-smoothing method is stable in calibration data and the choice of computing the smooth-factor.
5.4.1 Calibration data source
Figure 21 presents an accuracy comparison across different calibration data sources. We tested 8 samples from the GPQA dataset, 32 samples from the WG dataset, and 32 samples from the PQ dataset. The results indicate that the choice of calibration data has minimal impact on accuracy; all three datasets enable the W8A8 per-channel quantized model to achieve accuracy comparable to that of the FP16 counterpart.
5.4.2 Calibration data size
We evaluated the impact of calibration data volume on smooth-factor computation, comparing results from 8 samples versus a single sample from the GPQA dataset. Figure 22 demonstrates that the difference in performance is negligible; both resulting W8A8 models achieve accuracy comparable to their FP16 counterpart.
5.4.3 Smooth-factor calculation
Note that in Equation (1), we use “median” to calculate the smooth-factor. A natural alternative is to replace “median” by “maximum” in the following equation.
Figure 23 illustrates the comparison, revealing that there is negligible difference between using the “median” or “max” to calculate the smooth factor. This finding underscores the robustness of the bi-smoothing method in recovering the accuracy of W8A8 models.
5.4.4 Applying bi-smoothing to a subset of Transformer layers
Our previous analysis revealed that significant weight outliers are predominantly confined to the initial Transformer blocks. An ablation study on the number of blocks with bi-smoothing method was conducted, with results illustrated in Figure 24. Notably, applying bi-smoothing to only the 0th, 1st, and 3rd Transformer blocks yields W8A8 accuracy nearly equivalent to that of the full-precision (FP16) model. This finding corroborates our hypothesis that the initial blocks are the primary source of the LLaMA3-70B model series’ susceptibility to W8A8 quantization.
@@ 7.1 Post-training quantization
PTQ does not use any back-propagation to finetune weights. For weight-only quantization, LLM.int8() Dettmers et al. (2022a) uses a mixed precision with 16-bit MAC for outliers. GPTQ Frantar et al. (2023) and AWQ Lin et al. (2024a) quantize the per-group weight-only tensor using as small as 3-4 bits. SqueezeLLM Kim et al. (2024) uses non-uniform quantization and store outliers efficiently. For weight-activation quantization that enables faster inference Dettmers et al. (2022b); Wei et al. (2023) SmoothQuant Xiao et al. (2024) transfers the outliers from activation to weights to enable per-tensor W8A8. QServe Lin et al. (2024b) proposes W4A8K4 to serve LLM models. Some other studies enable even smaller number of bits Shao et al. (2024); Zhao et al. (2024); Ashkboos et al. (2024). Notably, our observations on the weights and activations of the LLaMA3-70B model series contrast with those of SmoothQuant. The proposed bi-smoothing method extends this approach by addressing both weight and activation outliers. Compared to other PTQ methods, we aim to maintain per-channel instead of per-group quantization for maximal hardware efficiency.
@@ 9.1 Models
| LLaMA3.1-70B-Instruct | https://huggingface.co/meta-llama/Llama-3.1-70B-Instruct |
@@
| LLaMA3.2-1B-Instruct | https://huggingface.co/meta-llama/Llama-3.2-1B-Instruct |
| LLaMA3.2-3B-Instruct | https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct |
@@
| Mistral-Large-Instruct-123B | https://huggingface.co/mistralai/Mistral-Large-Instruct-2407 |
@@ 9.2.1 The first block has more large outliers
We show the distribution of weights in the first block in LLaMA3-70B. Figures 25, 26, 27, 28, 29, 30, 31 shows seven weight matrices in the 0-th Transformer block. The z-axis shows the magnitudes of outliers in these matrices. We can see that the Q, K, V, Up, and Gate matrices have larger outliers and they form a “wall” in some input indices. The other two weights, namely O and Down, do not show such a pattern.